A bullet of mass m=50gm strikes t~0 a sand bag M=5kg hanging from a fixed point with a horizontal velocity vp if bullet sticks to the sand bag then find the final and initial ratio of kinetic energy

Grade:11

1 Answers

Sanju
6 years ago
I assume that you need the final and initial ratio of kinetic energy of bullet.
 
 According to Conservation of Momentum, mvp = (m+M)V where, m is the mass of the bullet, vp is the velocity of the bullet before collision, M is the mass of the sandbag and V is the velocity after the collision.            0.05* v= (0.050+5)VTherefore, vp = [5.050/0.05] V = 101 V Ratio of final KE to initial KE = [0.5mV2] / [0.5mvp2] = 10‑4  (approx.)